OpenStudy (anonymous):

What is the missing side length of an irregular pentagon with side lengths nineteen point nine centimeters, twenty-four point eight centimeters, thirty centimeters and forty-nine point two centimeters, if the perimeter is equal to the perimeter of a regular hexagon with side length twenty-six centimeters? A.six point one centimeters B.twenty-three point two centimeters C.thirty-two point one centimeters D.fifty-eight point one centimeters

OpenStudy (anonymous):

what the question to hard ? or to long ?

OpenStudy (secret66):

Pentagon=five sides. 19.9 + 24.8 + 30 + 49.2 = 123.9 cm 156 - 123.9 = 32.1 cm (final answer)

OpenStudy (a_clan):

perimeter of the regular hexagon= 26 * 6 perimeter of the irregular pentagon= 19.9 + 24.8 + 30 + 49.2 + x Since both the perimeter have to be equal, hence 26 * 6 = 19.9 + 24.8 + 30 + 49.2 + x